A wheel initially at 5 cycles per second decelerates and stops after completed 60 turns. What is the angular speed the wheel after completing 45 turns ?


1
Expert's answer
2021-11-29T11:42:21-0500
"w^2=2\alpha \theta\\5^2=2(60)\alpha\\\alpha=0.208\frac{rev}{s^2}\\5^2-w_f^2=2(45)(0.208)\\w=2.5\frac{rev}{s}"